# bayesOptimalControl.bib

### v1

@@ 1. Introduction
In optimal control, we aim to determine a control strategy for a dynamical system that minimizes a given cumulative running cost function over time, plus a terminal cost, within a finite time horizon. As general references, we refer to books [FlemingSoner1993] (providing a comprehensive introduction to the theory of SOC, focusing on Hamiltonian systems and Hamilton-Jacobi-Bellman (HJB) equations) and [BardiCapuzzoDolcetta1997] (exploring the connection between Optimal Control and the viscosity solutions of HJB equations).
In many applications which motivate our work, the system’s state evolves continuously over time, but the current state of the controlled system is not perfectly known. This discrepancy between the continuous evolution of the state and imperfect observations presents a significant challenge for control strategies. The books [Bensoussan1992, Mao2006, Krishnamurthy16] address control problems where the system state is only partially observable, using a filtering framework and applying it to problems of control under uncertainty.
@@
We start by describing the main frameworks for analyzing classical SOC problems. In the 1960s, Ronald A. Howard [Howard1960] popularized the term “Markov Decision Processes” (MDPs) and developed the policy iteration method, which is a fundamental technique for solving MDPs. Throughout the 1970s and 1980s, MDPs addressed more complex settings, including continuous-time processes and infinite-horizon problems, leading to the development of methods like value iteration and Q-learning. The integration of MDPs with machine learning, particularly reinforcement learning (RL), in the late 1980s and 1990s, see [BertsekasTsitsiklis1996, Szepesvari2010, sutton_barto18], among others, has led to significant advancements. This includes the development of algorithms like Temporal Difference (TD) learning and the popularization of Q-learning and deep reinforcement learning. Today, MDPs are central to many applications in artificial intelligence, robotics, economics, and operations research. The Kalman filter [Kalman1960] became an essential tool for dealing with linear systems with Gaussian noise, while the need for handling non-linear systems led to the development of the Extended Kalman Filter [WelchBishop1995], Ensemble Kalman Filter [Evensen1994] and, later, the Particle Filter in the 1990s [DoucetFreitasGordon2001]. The concept of separation principle in control theory, which suggests that control and filtering can be separated in some cases, was a significant result for linear systems but proved challenging in non-linear settings, see [AndersonMoore1979]. This book develops the separation principle in the context of linear systems and discusses its extension to non-linear systems through various filtering techniques.
The main alternative approach is based on the Hamilton-Jacobi-Bellman (HJB) equation. It describes the optimal cost function’s evolution in dynamic programming terms. This approach was extensively developed during the 1970s and 1980s. The application of viscosity solutions to HJB equations, cf. [Lions1982, CrandallLions1983, crandall@1984] by Michael Crandall and Pierre-Louis Lions in the 1980s provided critical mathematical tools for dealing with the challenges posed by the non-linearity and high dimensionality of realistic control problems. Modern research focuses on bridging the gap between theoretical optimality and practical computability, especially in high-dimensional spaces where traditional methods are computationally infeasible. Applications now span complex systems in finance, engineering, and networked systems, where uncertainty and partial observability are key concerns.
@@
An important technical tool required for deriving dynamic programming principles or Hamilton-Jacobi-Bellman equations for measure-valued stochastic processes is an appropriate Itô formula. There has been renewed interest in this problem, mainly coming from mean-field games and control of McKean–Vlasov equations. A very general Itô formula for measure-valued semi-martingales has recently been derived in [guo_pham_wei23] and references therein. More related works in the context of mean-field optimal control or optimal stopping, we refer to [talbi_touzi_zhang23a, talbi_touzi_zhang23b, germain_pham_warin22].
@@
The SOC problem with partial state observation is better understood when controls are relaxed i.e., $\alpha_{t}$ is replaced by a measure, see, for instance, [fleming_nisio84, elkaroui_du_jeanblanc88] for existence results for relaxed optimal controls. The classical work by Fleming [Fleming1980] introduced measure-valued processes for partially observed control problems, providing a theoretical foundation for analyzing stochastic control problems where the state is not fully observable. Fleming’s insights are essential for our analysis of measure-valued processes. Recent works have made significant strides in approximating SOC problems under partial observation. Tan and Yang [TanYang2023] discuss discrete-time approximation of continuous-time stochastic control problems under continuous time, partial observation. Their methods’ convergence properties support our work’s theoretical foundation, especially in the context of the measure-valued control framework we discuss. This recent work deals with approximating schemes for filtered problems, which may be relevant to our problem, which is not a standard filtering-control problem and only uses discrete time, partial observations.
@@
The situation changes quite drastically when we consider SOC problems formulated within the field of MDPs. Indeed, there is a classical and rich literature on so-called partially observed Markov decision processes (POMDPs), which, as the name suggests, is concerned with MDPs where only partial, noisy observation of the controlled state process are available. We refer to [Krishnamurthy16] for a monograph and [baeuerle_rieder17] for an interesting recent application to an economics problem. As discussed earlier, the strategy is to lift the POMDP by considering the conditional distribution of the unobserved full state. The resulting control problem for measure-valued (hence, in general, infinite-dimensional) processes is the seen to be a standard MDP, and can be analysed by standard methods. Note that, in contrast to the stochastic optimal control literature, the MDP literature is almost exclusively concerned with discrete time problems, and this also extends to the partially-observed case.
@@
The problem of SOC under noisy observation is also related to the problem of reinforcement learning [sutton_barto18]. Indeed, reinforcement learning operates under even less information, since not even knowledge of the driving dynamics of the controlled system is given, but rather has to be learned while controlling the system. Of course, stationarity of the system is usually assumed. Reinforcement learning is, again, generically formulated in discrete time, even though some attempts of continuous time extensions have recently been made, see, e.g., [wang_etal_20].
@@
The above setup invites to include a second control into our problem. As already indicated above, measurements are usually noisy (say, with standard deviation $\epsilon$), and may not convey information about the full state $X_{t}$, anyway, – think of noisy observations of a function of $X_{t}$, e.g., just a single component. However, in many cases, more precise measurement methods may be available, albeit at a higher cost. Hence, in addition to the control $\alpha$ guiding the state process $X_{t}$, we may consider a second control $\beta$ for the measurement method, as well as an associated cost term. In the statistical literature, this is also known as optimal experimental design, see, for instance, [smucker2018optimal]. Mathematically, this means that we obtain Bayesian updates $\mu_{t_{i}}=K_{\beta_{t_{i}}}(\mu_{t_{i}^{-}},Y_{t_{i}})\mu_{t_{i}^{-}}$ which are directly influenced by the control, not only indirectly via $\mu$ and $Y$. Of course, this principle may also be applied to the observation times themselves, which could, more generally, be chosen by the controller.
@@ 2. A motivating example
[…]
@@ 3.1. Formulation of the problem
To fix notation, let us first look at the dynamics of the uncontrolled process. To start out, we fix observation times $0<t_{1}<\cdots<t_{n}<T$, which are (for simplicity) assumed to be deterministic. Likewise, for simplicity we assume that there are no observations at times 00 and $T$, and we introduce $t_{0}\coloneqq 0$, $t_{n+1}\coloneqq T$. We also introduce the notation $\left\lfloor t\right\rfloor\coloneqq\max\Set{i}{t_{i}<t}$, with the convention that $\left\lfloor 0\right\rfloor\coloneqq 0$. Let $\mathcal{G}$ denote the infinitesimal generator of a $d$-dimensional diffusion process defined for $f\in C_{b}^{2}(\mathbb{R}^{d})$ by
@@
As such, it is easy to see that $\mu_{t}$ satisfies an "Itô-formula”. Given $\Phi:\mathcal{P}(\mathbb{R}^{d})\to\mathbb{R}$ “nice enough”, we associate to it its flat derivative, see, for instance, [dawson1993measure, Page 18f] or [kolokoltsov2010nonlinear, Appendix F].
@@
Assume $(\mu_{t})_{t\in[0,T]}$ solves (3.2) and that the function $\Phi(t,\mu)$ from $[0,T]\times\mathcal{P}(\mathbb{R}^{d})$ to $\mathbb{R}$ is differentiable w.r.t. the time variable $t$ and is in $\mathcal{S}^{1,1}(\mathcal{P}(\mathbb{R}^{d}))$ w.r.t. $\mu$.
@@
Sketch of proof, following [guo_pham_wei23]..
@@ 3.3. The HJB equation
(3.15) $u^{\ast}(t,\mu)\in\begin{cases}\operatorname*{arg\,min}_{(\alpha,\beta)\in U}\left\{\left\langle\mathcal{G}^{\ast}(\alpha)\mu\,,\frac{\delta V}{\delta\mu}(t,\mu,\cdot)\right\rangle+\ell(t,\mu,\alpha)\right\},&t\notin\set{t_{1},\ldots,t_{n}},\\ \operatorname*{arg\,min}_{(\alpha,\beta)\in U}\left\{h(t_{i},\mu,\beta)+E\left[V(t_{i},K_{\beta}(\cdot;\mu,Y^{\beta,\mu}_{i})\mu)\right]\right\},&t=t_{i}^{-},\ i\in\set{1,\ldots,n},\end{cases}$
@@
First we note that at observation times $t_{i}$, (3.12) is exactly (3.14b). We only need to derive (3.14a). Consider in (3.11) a constant control $u:=a=(\bar{\alpha},\bar{\beta})$ for some arbitrary $a\in U$. For any $t,\theta>0$ such that, for some $i=1,\ldots,n$, $[t,t+\theta]\subset[t_{i},t_{i+1})$, we have
[…]
@@
By the mean-value theorem, dividing by $\theta$ and then sending it to 00, yields
@@
$\widehat{u}(t,\mu)\in\begin{cases}\operatorname*{arg\,min}_{(\alpha,\beta)\in U}\left\{\left\langle\mathcal{G}^{\ast}(\alpha)\mu\,,p\right\rangle+\ell(t,\mu,\alpha)\right\},&t\notin\set{t_{1},\ldots,t_{n}},\\ \operatorname*{arg\,min}_{(\alpha,\beta)\in U}\left\{h(t_{i},\mu,\beta)+E\left[W(t_{i},K_{\beta}(\cdot;\mu,Y^{\beta,\mu}_{i})\mu)\right]\right\},&t=t_{i}^{-},\ i\in\set{1,\ldots,n}.\end{cases}$
@@
The HJB equation (3.14) is a piecewise backward PDE. Given the solution $V(t_{i}^{-},\cdot)$ at time $t_{i}$, $i=1,\ldots,n+1$ (interpreted as $V(t_{n+1}^{-},\mu)=g(\mu)$ at $t_{n+1}=T$), we obtain the solution on $[t_{i-1},t_{i})$ by solving the PDE (3.14a) backward in time. Then, we define $V(t_{i-1}^{-},\mu)$ by (3.14b), and continue as before.
@@
Suppose that we are given a subset $\mathcal{M}\subset\mathcal{P}(\mathbb{R}^{d})$, which is invariant under $\mathcal{G}^{\ast}(\alpha)$ as well as under the Bayesian update for any optimal control $u$, and that we start in $\mu_{0}\in\mathcal{M}$. Let us further assume that probability measures in $\mathcal{M}$ are uniquely characterized by the expectations of functions $\Set{\varphi_{j}}{j\in\mathcal{J}}$ indexed by a (finite or infinite) index set $\mathcal{J}$.
@@ 4. Optimal control of an Ornstein – Uhlenbeck process
It is well–known, see [guo_pham_wei23], that the value function to the stochastic optimal control problem – with no observation of the trajectories – is given as
@@
$\quad\quad+\int_{\mathbb{R}^{d}}x^{2}\mu(\mathrm{d}x)+C\alpha^{2}\biggr{\}}$
@@
$=\inf_{\alpha\in\mathbb{R}}\biggl{\{}C\alpha^{2}+\left(\frac{\partial U}{\partial m}(t,m,s)+2m\frac{\partial U}{\partial s}(t,m,s)\right)u+$
@@ 5. Numerical approach
5.1. Choice of the computational domain and boundary conditions
As mentioned, the original problem is defined on an unbounded domain with $m\in(-\infty,\infty)$ and $z\in[0,\infty)$. However, to solve the equation numerically, one must choose the truncated domain. Furthermore, one should either prescribe the boundary conditions at the boundaries of the chosen domain, or propose the numerical scheme, which uses only the interior points to compute current values and is shown to be stable. The upwind difference scheme has the desired property as long as information propagates from the initial conditions by lines flowing outside the domain. To check if this is the case, one can compute the characteristic lines of the equation following section 3.2 of [evans@2010]. In our case, the characteristic lines of (5.1a) are flowing from outside the domain, which is unwanted. However, if we restate the problem in a reversed time flow as follows
@@ 5.1. Choice of the computational domain and boundary conditions
we can choose a domain, such that the characteristic lines cross the boundary from inside the domain. See, for example, Figure 1, which shows the characteristic lines for $m\in[-1,1]$ and $z\in[0,1]$ for the synthetic problem with parameters given in subsection 5.5. The lines for $m$ are symmetric around $m(0)=0$; for $z$ they are symmetric around $z=b^{2}/(2\theta)=0.5$. Therefore, as long as the computational domain contains $m=0$ and $z=0.5$, the numerical solution is stable. The details on the derivation of the characteristics are presented in Appendix A. Additionally, Saldi, Basar, and Raginsky [SaldiBasarRaginsky2018] explored the asymptotic optimality of finite model approximations for partially observed Markov decision processes (POMDPs). Although relevant for a time discrete setting, their approximation results for POMDPs may provide theoretical support for our domain truncation numerical approach as well, connecting our work to established results in the approximation of partially observed control problems.
@@
5.2. Numerical scheme for HJB equation
In order to construct a convergent scheme we will follow the guidelines outlined in section 3.1 of [falcone@2016] and [crandall@1984]. We introduce the grid

### v2

# Continuous time Stochastic optimal control under discrete time partial observations

@@ 1. Introduction
In optimal control, we aim to determine a control strategy for a dynamical system that minimizes a given cumulative running cost function over time, plus a terminal cost, within a finite time horizon. As general references, we refer to books [FS93] (providing a comprehensive introduction to the theory of SOC, focusing on Hamiltonian systems and Hamilton-Jacobi-Bellman (HJB) equations) and [BC97] (exploring the connection between Optimal Control and the viscosity solutions of HJB equations).
In many applications which motivate our work, the system’s state evolves continuously over time, but the current state of the controlled system is not perfectly known. This discrepancy between the continuous evolution of the state and imperfect observations presents a significant challenge for control strategies. The books [Ben92, Mao06, Kri16] address control problems where the system state is only partially observable, using a filtering framework and applying it to problems of control under uncertainty.
@@
We start by describing the main frameworks for analyzing classical SOC problems. In the 1960s, Ronald A. Howard [How60] popularized the term “Markov Decision Processes” (MDPs) and developed the policy iteration method, which is a fundamental technique for solving MDPs. Throughout the 1970s and 1980s, MDPs addressed more complex settings, including continuous-time processes and infinite-horizon problems, leading to the development of methods like value iteration and Q-learning. The integration of MDPs with machine learning, particularly reinforcement learning (RL), in the late 1980s and 1990s, see [BT96, Sze10, SB18], among others, has led to significant advancements. This includes the development of algorithms like Temporal Difference (TD) learning and the popularization of Q-learning and deep reinforcement learning. Today, MDPs are central to many applications in artificial intelligence, robotics, economics, and operations research. The Kalman filter [Kal60] became an essential tool for dealing with linear systems with Gaussian noise, while the need for handling non-linear systems led to the development of the Extended Kalman Filter [WB95], Ensemble Kalman Filter [Eve94] and, later, the Particle Filter in the 1990s [DFG01]. The concept of separation principle in control theory, which suggests that control and filtering can be separated in some cases, was a significant result for linear systems but proved challenging in non-linear settings, see [AM79]. This book develops the separation principle in the context of linear systems and discusses its extension to non-linear systems through various filtering techniques.
The main alternative approach is based on the Hamilton-Jacobi-Bellman (HJB) equation. It describes the optimal cost function’s evolution in dynamic programming terms. This approach was extensively developed during the 1970s and 1980s. The application of viscosity solutions to HJB equations, cf. [Lio82, CL83, CL84] by Michael Crandall and Pierre-Louis Lions in the 1980s provided critical mathematical tools for dealing with the challenges posed by the non-linearity and high dimensionality of realistic control problems. Modern research focuses on bridging the gap between theoretical optimality and practical computability, especially in high-dimensional spaces where traditional methods are computationally infeasible. Applications now span complex systems in finance, engineering, and networked systems, where uncertainty and partial observability are key concerns.
@@
An important technical tool required for deriving dynamic programming principles or Hamilton-Jacobi-Bellman equations for measure-valued stochastic processes is an appropriate Itô formula. There has been renewed interest in this problem, mainly coming from mean-field games and control of McKean–Vlasov equations. A very general Itô formula for measure-valued semi-martingales has recently been derived in [GPW23] and references therein. More related works in the context of mean-field optimal control or optimal stopping, we refer to [TTZ23, TTZ23a, GPW22].
@@
The SOC problem with partial state observation is better understood when controls are relaxed i.e., $\alpha_{t}$ is replaced by a measure, see, for instance, [FN84, ENJ88] for existence results for relaxed optimal controls. The classical work by Fleming [Fle80] introduced measure-valued processes for partially observed control problems, providing a theoretical foundation for analyzing stochastic control problems where the state is not fully observable. Fleming’s insights are essential for our analysis of measure-valued processes. Recent works have made significant strides in approximating SOC problems under partial observation. Tan and Yang [TY23] discuss discrete-time approximation of continuous-time stochastic control problems under continuous time, partial observation. Their methods’ convergence properties support our work’s theoretical foundation, especially in the context of the measure-valued control framework we discuss. This recent work deals with approximating schemes for filtered problems, which may be relevant to our problem, which is not a standard filtering-control problem and only uses discrete time, partial observations.
@@
The situation changes quite drastically when we consider SOC problems formulated within the field of MDPs. Indeed, there is a classical and rich literature on so-called partially observed Markov decision processes (POMDPs), which, as the name suggests, is concerned with MDPs where only partial, noisy observation of the controlled state process are available. We refer to [Kri16] for a monograph and [BR17] for an interesting recent application to an economics problem. As discussed earlier, the strategy is to lift the POMDP by considering the conditional distribution of the unobserved full state. The resulting control problem for measure-valued (hence, in general, infinite-dimensional) processes is the seen to be a standard MDP, and can be analysed by standard methods. Note that, in contrast to the stochastic optimal control literature, the MDP literature is almost exclusively concerned with discrete time problems, and this also extends to the partially-observed case.
@@
The problem of SOC under noisy observation is also related to the problem of reinforcement learning [SB18]. Indeed, reinforcement learning operates under even less information, since not even knowledge of the driving dynamics of the controlled system is given, but rather has to be learned while controlling the system. Of course, stationarity of the system is usually assumed. Reinforcement learning is, again, generically formulated in discrete time, even though some attempts of continuous time extensions have recently been made, see, e.g., [WZZ20].
@@
The above setup invites to include a second control into our problem. As already indicated above, measurements are usually noisy (say, with standard deviation $\epsilon$), and may not convey information about the full state $X_{t}$, anyway, – think of noisy observations of a function of $X_{t}$, e.g., just a single component. However, in many cases, more precise measurement methods may be available, albeit at a higher cost. Hence, in addition to the control $\alpha$ guiding the state process $X_{t}$, we may consider a second control $\beta$ for the measurement method, as well as an associated cost term. In the statistical literature, this is also known as optimal experimental design, see, for instance, [SKA18]. Mathematically, this means that we obtain Bayesian updates $\mu_{t_{i}}=K_{\beta_{t_{i}}}(\mu_{t_{i}^{-}},Y_{t_{i}})\mu_{t_{i}^{-}}$ which are directly influenced by the control, not only indirectly via $\mu$ and $Y$. Of course, this principle may also be applied to the observation times themselves, which could, more generally, be chosen by the controller.
@@ 2. A motivating example
(2.2) $J(\alpha)\coloneqq\mathbb{E}\left[\int_{0}^{T}\left(\int_{\mathbb{R}}x^{2}\mu^{\alpha}_{t}(\mathrm{d}x)+C\alpha_{t}^{2}\right)dt+\int_{\mathbb{R}}x^{2}\mu^{\alpha}_{T}(\mathrm{d}x)+\sum_{i=1}^{n}\frac{1}{\beta_{i}}\right].$
@@ 3.1. Formulation of the problem
[…]
@@
As such, it is easy to see that $\mu_{t}$ satisfies an "Itô-formula”. Given $\Phi:\mathcal{P}(\mathbb{R}^{d})\to\mathbb{R}$ “nice enough”, we associate to it its flat derivative, see, for instance, [Daw93, Page 18f] or [Kol10, Appendix F].
@@
Assume $(\mu_{t})_{t\in[0,T]}$ solves (3.2) and that the function $\Phi(t,\mu)$ from $0,T]\times\mathcal{P}(\mathbb{R}^{d})$ to $\mathbb{R}$ is differentiable w.r.t. the time variable $t$ and is in $\mathcal{S}^{1,1}(\mathcal{P}(\mathbb{R}^{d}))$ w.r.t. $\mu$.
@@
Sketch of proof, following [GPW23]..
@@ 3.3. The HJB equation
[…]
@@
First we note that at observation times $t_{i}$, (3.12) is exactly (3.14b). We only need to derive (3.14a). Consider in (3.11) a constant control $u:=a=(\bar{\alpha},\bar{\beta})$ for some arbitrary $a\in U$. For any $t,\theta>0$ such that, for some $i=1,\ldots,n$, $t,t+\theta]\subset[t_{i},t_{i+1})$, we have
$\begin{array}[]{ll}V(t,\mu)\leq\mathbb{E}\left[\int_{t}^{t+\theta}\ell(s,\mu_{s}^{a}(t,\mu),a)\,ds+V(t+\theta,\mu^{a}_{t+\theta}(t,\mu))\right].\end{array}$
@@
By the mean-value theorem, dividing by $\theta$ and then sending it to 0, yields
@@
[…]
@@
The HJB equation (3.14) is a piecewise backward PDE. Given the solution $V(t_{i}^{-},\cdot)$ at time $t_{i}$, $i=1,\ldots,n+1$ (interpreted as $V(t_{n+1}^{-},\mu)=g(\mu)$ at $t_{n+1}=T$), we obtain the solution on $t_{i-1},t_{i})$ by solving the PDE (3.14a) backward in time. Then, we define $V(t_{i-1}^{-},\mu)$ by (3.14b), and continue as before.
@@
[…]
@@ 4. Optimal control of an Ornstein – Uhlenbeck process
It is well–known, see [GPW23], that the value function to the stochastic optimal control problem – with no observation of the trajectories – is given as
@@
$\quad\quad+\int_{\mathbb{R}}x^{2}\mu(\mathrm{d}x)+C\alpha^{2}\biggr{\}}$
@@
$=\inf_{\alpha\in\mathbb{R}}\biggl{\{}C\alpha^{2}+\left(\frac{\partial U}{\partial m}(t,m,s)+2m\frac{\partial U}{\partial s}(t,m,s)\right)\alpha+$
@@ 5. Numerical approach
We illustrate our method on three numerical examples: a classic linear–quadratic control problem, control of measurement noise, and control under a high-cost region constraint.
5.1. Example 1: Linear-quadratic control under uncertain observation
@@
5.1.1. Choice of the computational domain and boundary conditions
As mentioned, the original problem is defined on an unbounded domain with $m\in(-\infty,\infty)$ and $z\in[0,\infty)$. However, to solve the equation numerically, one must choose the truncated domain. Furthermore, one should either prescribe the boundary conditions at the boundaries of the chosen domain, or propose the numerical scheme, which uses only the interior points to compute current values and is shown to be stable. The upwind difference scheme has the desired property as long as information propagates from the initial conditions by lines flowing outside the domain. To check if this is the case, one can compute the characteristic lines of the equation following section 3.2 of [Eva10]. In our case, the characteristic lines of (5.38a) are flowing from outside the domain, which is unwanted. However, if we restate the problem in a reversed time flow as follows
@@ 5.1.1. Choice of the computational domain and boundary conditions
we can choose a domain, such that the characteristic lines cross the boundary from inside the domain. See, for example, Figure 1, which shows the characteristic lines for $m\in[-1,1]$ and $z\in[0,1]$ for the synthetic problem with parameters given in subsection 5.1.5. The lines for $m$ are symmetric around $m(0)=0$; for $z$ they are symmetric around $z=b^{2}/(2\theta)$. Therefore, as long as the computational domain contains $m=0$ and $z=0.5$, the numerical solution is stable. The details on the derivation of the characteristics are presented in Appendix A. Additionally, Saldi, Basar, and Raginsky [SBR18] explored the asymptotic optimality of finite model approximations for partially observed Markov decision processes (POMDPs). Although relevant for a time discrete setting, their approximation results for POMDPs may provide theoretical support for our domain truncation numerical approach as well, connecting our work to established results in the approximation of partially observed control problems.
@@
5.1.2. Numerical scheme for HJB equation
In order to construct a convergent scheme we will follow the guidelines outlined in section 3.1 of [FF16] and [CL84]. We introduce the grid